### v1

# Sequential Multi-Object Grasping with One Dexterous Hand

Abstract
Sequentially grasping multiple objects with multi-fingered hands is common in daily life, where humans can fully leverage the dexterity of their hands to enclose multiple objects. However, the diversity of object geometries and the complex contact interactions required for high-DOF hands to grasp one object while enclosing another make sequential multi-object grasping challenging for robots. In this paper, we propose SeqMultiGrasp, a system for sequentially grasping objects with a four-fingered Allegro Hand. We focus on sequentially grasping two objects, ensuring that the hand fully encloses one object before lifting it and then grasps the second object without dropping the first. Our system first synthesizes single-object grasp candidates, where each grasp is constrained to use only a subset of the hand’s links. These grasps are then validated in a physics simulator to ensure stability and feasibility. Next, we merge the validated single-object grasp poses to construct multi-object grasp configurations. For real-world deployment, we train a diffusion model conditioned on point clouds to propose grasp poses, followed by a heuristic-based execution strategy. We test our system using $8\times 8$ object combinations in simulation and $6\times 3$ object combinations in real. Our diffusion-based grasp model obtains an average success rate of 65.8% over 1600 simulation trials and 56.7% over 90 real-world trials, suggesting that it is a promising approach for sequential multi-object grasping with multi-fingered hands. Supplementary material is available on our project website: https://hesic73.github.io/SeqMultiGrasp.
I Introduction
Despite significant advancements and interest in multi-fingered dexterous hand manipulation [1, 2, 3, 4, 5, 6], much of the research in robotic grasping with such hands remains centered on single-object scenarios. This often reduces the sophisticated capabilities of dexterous hands to those of simpler grippers, thereby under-utilizing their potential. The high degrees of freedom of human hands, however, enable simultaneous manipulation and independent control of multiple objects. This capability is fundamental for complex tasks such as multi-object grasping [7], in-hand manipulation [8, 1], and object separation [9].
Recent efforts in dexterous grasping have leveraged dexterous hands for multi-object manipulation. For example, MultiGrasp [7] presents a framework for simultaneously grasping multiple objects using a Shadow Hand, and focuses on generating pre-grasp poses and executing lifting strategies. However, MultiGrasp targets scenarios where objects are positioned in close proximity, limiting its applicability and overlooking the potential of dexterous hands for independent control of multiple objects. Similarly, [10] explores multi-object grasping but assumes a fixed hand pose and requires human intervention to place objects within the hand. In contrast, we address the relatively unexplored domain of autonomous and sequential multi-object grasping, where a robot must secure one object, and then grasp the second object while maintaining control of the first.
Sequentially grasping multiple objects is particularly beneficial when objects are located too far apart for simultaneous manipulation. While one option might be to push objects together [11, 12] as a preprocessing step before simultaneous grasping, this might not always be feasible with high-friction surfaces and risks toppling certain objects.
Our method, SeqMultiGrasp, builds on the Differentiable Force Closure (DFC) algorithm [13] to generate grasp poses that utilize a subset of the hand links to secure one object. These grasps are then evaluated in a physics simulator, where unstable or infeasible configurations are filtered out. The validated grasp poses are subsequently merged to construct stable configurations for sequentially grasping two objects. Leveraging this dataset, we train a generative model conditioned on point cloud inputs to propose grasp poses. Finally, we employ motion planning and heuristic strategies to execute precise grasping actions, enabling real-world multi-object manipulation.
Figure 1: Overview of SeqMultiGrasp in action using a Franka Panda robot arm and a four-fingered Allegro Hand. From an initial scene with two objects (1), our method first enables grasping and lifting the first object (a fake plum) with two fingers (2). Then, the hand rotates so that it can grasp (3) and lift (4) the can, without losing control of the plum.
The contributions of this work are: (i) a novel framework, SeqMultiGrasp, for real-world sequential multi-object grasping; (ii) a new open-source dataset specifically designed to train sequential multi-object grasping methods; (iii) the first real-world robotic experiments demonstrating sequential multi-object grasping, where we conduct a total of 180 trials across 18 object pairs, with 90 trials replaying grasps from our dataset and 90 trials using grasp poses generated by our diffusion model.
II Related Work
II-A Dexterous Grasping
Robotic grasping is one of the most fundamental tasks in robotics and is often a prerequisite for subsequent manipulation tasks. Research in this area can be broadly categorized as analytic versus data-driven. In analytic methods, it is standard to use geometric and mathematical optimization to determine suitable grasps [14] that can form force closure on an object and resist external forces [15, 16]. For example, [17, 18] analyze force-closure grasps for grippers with multiple fingers. A recent work, FRoGGeR [19] uses a differentiable min-weight metric for fast grasp refinement. While impressive, analytic methods typically assume knowledge of object geometries and contact locations.
Data-driven methods for grasping [20, 21] use machine learning methods to learn suitable end-effector poses from data. One direction has been with learning for parallel-jaw grasping [22] such as in Dex-Net [23] and AnyGrasp [24]. However, using a parallel-jaw gripper may restrict flexibility in grasping certain objects. Thus, researchers have built large-scale datasets of grasping with multi-fingered hands. As an initial step, [25] generate 6.9K grasps using the GraspIt! [26] software tool. Other data-generation procedures such as Grasp’D [27] and Fast-Grasp’D [28] use differentiable contact simulation to generate diverse grasps, with TaskDexGrasp [29] additionally focusing on task-oriented dexterous hand pose synthesis for both prehensile and non-prehensile tasks. Another tactic is to instead use human demonstrations and retarget them to robot hands [30].
In highly relevant work, DexGraspNet [3] contains 1.32M grasps by using a modified Differentiable Force Closure (DFC) [13] method, and has improved grasp quality compared to other datasets [31]. Likewise, methods such as [2, 4] generate large-scale datasets by synthesizing grasps in simulation, where grasp poses are initialized, optimized using an energy-based formulation, and evaluated in a physics simulator. We extend this pipeline to support sequentially grasping multiple objects.
II-B Multi-Object Grasping
While most robotic grasping research studies grasping single objects, grasping multiple objects is an important research direction. As with single object grasping, multi-object grasping research can be broadly divided into analytic and data-driven grasping procedures, and furthermore, into grasping with simple (e.g., parallel-jaw) or complex grippers.
Early work in multi-object grasping focused on kinematic analysis and internal force modeling to characterize the relationship between the robot hand and multiple objects for grasp stability [32, 33, 34]. Some of these works also assume that the hand already encloses the objects in preparation for grasping [35, 36]. These works primarily focus on theoretical grasp analysis rather than practical grasp synthesis and execution. In addition, none of them study sequential grasping.
Among recent analytic methods, Agboh et al. [11] study grasping multiple objects on a plane with a parallel-jaw gripper and Yao et al. [10] develop an algorithm for multi-object grasping with a multi-fingered hand that utilizes pairwise contacts on arbitrary opposing hand surface regions. In contrast to [11], we use a higher-DOF hand which offers greater flexibility in grasping more diverse objects. Furthermore, while [10] show multi-object grasping, it requires humans to directly place the objects on the robot hand.
Another direction is to focus on the analysis and design specialized hardware systems for multi-object grasping [37]. Our focus is on expanding the capabilities of existing hands that are not specialized towards multi-object grasping.
An alternative to analytic methods for multi-object grasping is to use machine learning. One direction focuses on orthogonal tasks such as estimating the number of objects grasped from a container [38]. In more closely-related work, MultiGrasp [7] proposes a method to simultaneously grasp multiple objects with a high-DOF Shadow Hand. Their method uses a generative model with goal-conditioned reinforcement learning. MultiGrasp assumes that all objects are within close proximity to allow one grasp to succeed. However, this can be a strong assumption in practice, and moving objects closer to each other as a preprocessing procedure (via pushing or pick-and-place) may increase the risk of failures. Furthermore, experiments were mostly shown in simulation, whereas we show substantial real-world experiments.
III Problem Statement and Assumptions
We study sequential multi-object grasping on a tabletop surface with $N_{o}$ objects, denoted as $\mathbf{O}=\{O_{j}\}_{j=1}^{N_{o}}$.
To represent each object $O_{j}$, we use a point cloud $P_{j}$ sampled from its surface $S(O_{j})$. To describe a single object, we can use $O$ and suppress the subscript when the distinction is not needed. Our objective is to generate a sequence of grasping actions $\mathcal{A}=\{a^{1},a^{2},\dots,a^{L}\}$ such that a robot sequentially grasps and lifts multiple objects while maintaining grasp stability for all objects in hand. We assume a robot arm with a single multi-fingered hand. In this work, we use a Franka Panda arm equipped with an Allegro Hand (see Fig. 1). We focus on when $N_{o}=2$, i.e., grasping two objects sequentially, and where the objects are sufficiently far apart that they cannot be grasped simultaneously with one hand. A trial is an instance of this sequential grasping problem, and is considered successful if both objects are lifted off the tabletop in sequence and can be stabilized for 2 seconds.
IV Proposed Method
Figure 2: Overview of SeqMultiGrasp. (a) Contact candidates for different grasp types, where red and blue dots are used for pinch-like and side grasps, respectively (Sec. IV-A and IV-B). (b) Multi-object grasp configuration generation, where validated single-object grasps are merged into feasible multi-object grasps (Sec. IV-BIV-C, and IV-D). (c) Real-world grasp proposal process, where a diffusion model conditioned on point clouds generates grasp poses (Sec. IV-E). (d) Illustration of heuristic-based sequential grasping execution (Sec. IV-F).
SeqMultiGrasp consists of the following steps. First, we propose to divide up sequential grasping into two types of grasps (Section IV-A). Second, we generate a dataset of single-object grasp poses (Section IV-B) which we then validate in simulation (Section IV-C). Next, we merge multi-object grasps (Section IV-D). We use our data to train a diffusion model (conditioned on point clouds) to generate poses for grasp execution (Section IV-E). In Section IV-F, we discuss remaining implementation details. See Fig. 2 for an overview of SeqMultiGrasp.
IV-A Hand Pose Configuration and Grasp Styles
To achieve sequential grasping, our approach first proposes a unified hand configuration:
$H=(\theta,\{T_{j}\}_{j=1}^{N_{o}}),$ (1)
where $\theta\in\mathbb{R}^{d}$ represents the joint configuration of the robotic hand, and $T_{j}\in SE(3)$ denotes the relative pose of each object $O_{j}$ with respect to the hand. Given the proposed $H$, we generate the sequential grasping action sequence $\mathcal{A}$ to ensure all designated objects are grasped one by one.
As shown in Lum et al. [2], robust sim-to-real transfer for dexterous grasping requires multiple contact points between the robotic hand and the object. With this constraint in mind, we propose the following sequential grasping sequence:
  • First Grasp (Pinch-Like Grasp): The robot lifts the first object using any combination of the thumb, index, and middle fingers.
  • Second Grasp (Side grasp): The robot secures the second object using its ring finger and palm.
We empirically find that this grasping strategy can lead to viable and stable multi-object grasping given the Allegro Hand’s hardware. We leave exploration of other sequential dexterous grasping strategies to future work.
IV-B Dataset Preparation: Single-object Grasp Synthesis
We propose to construct a dataset of grasps for sequential multi-object grasping, so that we can use it as training data for a grasp synthesis model. Since no dataset exists for sequential multi-object grasping, we create our own. The closest is the grasping dataset from MultiGrasp [7], but it uses the Shadow Hand instead of the Allegro Hand, and is not designed for sequential grasping. To build our dataset, we first synthesize single-object grasp poses by building upon the DFC algorithm [13] which casts the grasp synthesis problem as minimizing an energy function via optimization.
In the single-object grasping case, the hand configuration simplifies to $H=(\theta,T)$, and $H$ follows a Gibbs distribution:
$p(H|\mathbf{O})=\frac{p(H,\mathbf{O})}{p(\mathbf{O})}\propto p(H,\mathbf{O})\sim\frac{1}{Z}e^{-E(H,\mathbf{O})},$ (2)
where $Z$ is the partition function and $E(H,\mathbf{O})$ represents the energy associated with the hand-object configuration.
Following [2, 3, 4], we use the energy function:
$E=E_{fc}+w_{dis}E_{dis}+w_{p}E_{p}+w_{sp}E_{sp}+w_{q}E_{q},$ (3)
where $E_{fc}$ is the force closure term [13], $E_{dis}$ penalizes the distance between contact points and the object’s surface, and $E_{p}$ penalizes penetration between the hand, object, and tabletop. The $E_{sp}$ term penalizes self-penetration of the hand, and $E_{q}$ penalizes joint limit violations. The $w$ terms are coefficients to weigh the components of $E$.
To generate single-object grasp configurations, we first sample contact points from contact candidates (see Fig. 2-a) on the hand surface and initialize the configurations. We then optimize them with a gradient-based approach combined with the Metropolis-Adjusted Langevin Algorithm (MALA). Finally, configurations with energy exceeding a predefined threshold are filtered out.
While our grasp synthesis pipeline builds upon existing work [2, 3, 4], we introduce several modifications to better support sequential multi-object grasping:
  • We restrict contact candidates to specific hand links for the first and second objects, based on our characterization of pinch-like and side grasps from Section IV-A.
  • The initialization method is designed to better suit the tabletop grasping scenario while favoring pinch-like and side grasps. For pinch-like grasps, we adopt the initialization method from [7]. For side grasps, we manually design canonical poses and introduce randomization.
  • Additional stability considerations are necessary since sequential multi-object grasping involves grasping a second object while the first remains held and undergoes motion. To address this, we adopt a hybrid grasp validation approach (see Section IV-C).
IV-C Grasp Validation in Simulation
While the optimization procedure from Section IV-B creates grasp candidates, we still require validation to ensure that they lead to successful grasps. Thus, we execute grasps in a GPU-accelerated physics simulator, ManiSkill [39], to validate them while filtering out unsuccessful candidates.
[…]
Rotation Robustness: From [3], we evaluate grasp stability under six different gravity directions. A grasp is considered valid if it maintains force closure across all tested conditions.
Execution Feasibility: Adapted from [2], we verify whether the grasp can be successfully executed without causing collisions with the environment. Our validation is simpler: we do not distinguish between different failure reasons, nor do we conduct multiple trials to smooth labels.
IV-D Merging Multi-object Grasps
After validating the stability and feasibility of single-object grasp poses, the next step is to combine them into multi-object grasp configurations (see Fig. 2-b). Multi-object grasp poses are generated by merging single-object grasps when their associated hand links and joints are completely disjoint. In single-object grasp synthesis, the DFC algorithm determines a hand configuration along with its contact points. Multi-object grasps can be formed by combining single-object grasps, provided that the contact points of different objects do not reside on the same finger. When merging, the joint angles of a finger are set according to the contact points of the object it is grasping. If a finger is not involved in grasping any object, its joint angles are inherited randomly from one of the single-object grasps. This merging process enables the simultaneous grasping of multiple objects while maintaining non-overlapping control constraints.
Figure 3: The objects we test in simulation (left) and real (right) for multi-object grasping. In both images, we use a dashed blue box to indicate the set of objects that the robot grasps first with the pinch grasp. Similarly, the dashed red box shows the set of objects that are grasped second with a side grasp. See Tables I and III for the name of these objects in simulation and real, respectively. The annotated numbers for simulation objects match those in Table I.
IV-E Diffusion-based Pose Generation
Our method for generating grasp poses provides a foundation for sequential multi-object grasping. However, generating these grasp poses for a new object pair for sequential grasping can be computationally expensive due to the optimization of the energy function. To explore alternative methods for generating grasp poses, we use a diffusion model [40] to generate hand poses conditioned on object point clouds $\mathbf{P}=\{P_{j}\}_{j=1}^{N_{o}}$ (see Fig. 2-c). Diffusion models are well-suited for this task as they can learn complex, high-dimensional distributions [41]. The forward process gradually corrupts the hand configuration $H$ with Gaussian noise:
$q(H_{t}|H_{t-1})=\mathcal{N}\left(H_{t};\sqrt{1-\beta_{t}}H_{t-1},\beta_{t}\mathbf{I}\right),$ (4)
where $\beta_{t}$ controls the noise level and $\mathbf{I}$ is the identity matrix. The reverse process reconstructs $H_{0}$ by iteratively denoising:
$p_{\phi}(H_{t-1}|H_{t},\mathbf{P})=\mathcal{N}\left(H_{t-1};\mu_{\phi}(H_{t},t,\mathbf{P}),\Sigma_{\phi}(H_{t},t,\mathbf{P})\right).$ (5)
Here, $\mu_{\phi}$ and $\Sigma_{\phi}$ are the predicted mean and covariance, respectively. The model is trained to predict the added noise and minimize the denoising error.
The network architecture and data representation build upon prior work [7, 41, 2, 42, 3]. We employ PointNet++ [43] for point cloud feature extraction and use a rotation matrix to represent object orientation relative to the hand, applying singular value decomposition (SVD) [44] to the network’s prediction to ensure orthogonality.
IV-F Other Implementation Details
We now discuss a grasp execution heuristic which we find to be important in practice. One approach to grasp execution [5, 6, 7] involves training reinforcement learning (RL) policies. However, these methods require complex reward design for training and often suffer from significant sim-to-real transfer challenges.
Instead, we adopt a heuristic-based approach inspired by prior work [4, 2], which avoids the complexities of RL. Specifically, we employ a simple squeeze-and-lift procedure. Given a hand configuration, we follow the strategy in Section IV-A to determine which fingers are used for each object. During grasp execution, we first use CuRobo [45] for motion planning to move the end effector to a collision-free pose set at a position offset from the grasp pose. Once in place, the end effector moves slowly toward the grasp pose without further collision checking.
[…]
This heuristic approach is simple and avoids the complexities of learning-based policies. However, it has limitations: it relies on predefined finger assignments and may cause unintended collisions, potentially leading to grasp failures. Despite these limitations, we find that this approach performs reliably in our multi-object grasping setup.
V Experiments
V-A Object Selection and Data Generation Evaluation
For experiments in simulation, we select 8 objects for pinch-like grasping and 8 objects for side-grasping from YCB [46], ContactDB [47] and Omni6DPose [48]. Some objects used for pinch-like grasping are rescaled to ensure they can be stably grasped by a few fingers. See Fig. 3 (left).
To evaluate our dataset generation pipeline, we report the synthesis success rates for pinch-grasping objects and side-grasping objects (Table I). Due to the use of fewer hand links and our strict hybrid validation criteria, the success rates are lower compared to single-object grasping scenarios [3, 4, 2]. However, since the entire process is parallelized, this approach remains feasible despite the lower raw success rates. We store all resulting successful grasps in our data for subsequent experiments.
Pinch Grasp Side Grasp
Objects Success Objects Success
0. plum 0.4% 0. slender cylinder 3.3%
1. cube 3.3% 1. medicine bottle 1 1.7%
2. duck 2.0% 2. soda can 2.1%
3. small cylinder 0.2% 3. soup can 4.1%
4. bunny 3.3% 4. boxed beverage 5.6%
5. lemon 0.1% 5. shaving foam 1.2%
6. train 1.0% 6. medicine bottle 2 1.9%
7. bottle 1.6% 7. canned meat 3.9%
TABLE I: The objects we test in ManiSkill simulation for the first (pinch) grasp and the second (side) grasp (see Fig. 3 (left)). We show the associated grasp synthesis success rate.
V-B Simulation Experiment Setup
[…]
We conduct the following experiments in simulation:
  • Dataset Replay (DR): We evaluate grasp poses sampled from our dataset. While each single-object grasp has already passed strict validation, failure may still occur after merging. This is mainly due to three factors: (1) Object placement shifts may lead to motion planning failures. (2) During execution, unintended interactions between the two objects or between the hand and objects may cause slippage or collisions. (3) A slight discrepancy exists between the validation and evaluation environments, as validation uses the PhysX GPU backend, while evaluation runs on the CPU.
  • Learned Grasp (LG): We sample grasp poses from our diffusion model conditioned on point clouds and execute them in simulation to measure the success rate.
  • Ablation studies: We compare rotation representations for network prediction, including Euler angles, axis-angle, quaternion, 6D representation [49], and our chosen rotation matrix (orthogonalized via SVD). We test each of these representations over all $8\times 8$ object combinations using the LG experiment protocol.
Figure 4: A heat map where each grid cell shows the success rate (from 25 trials) of sequentially grasping two objects in simulation for dataset replay experiments. The first grasp is a pinch grasp (vertical axis) and the second grasp is a side grasp (horizontal axis). We refer to objects by an integer index (0 through 7). Table I lists the object names for each index.
V-C Results from Simulation Experiments
Figure 5: Example trials of SeqMultiGrasp in ManiSkill simulation (top row) and in real world experiments (remaining rows). The first three images in each row show the hand moving towards, grasping, and then lifting the first object. The next three images show how the hand rotates, approaches, and then lifts the second object, which is cylindrical in shape. See Figure 3 for the objects that we test in simulation and the real world. All rollouts here are successful except for the last one, which shows a representative failure mode when grasping the second object, as the hand lost control of the Pepsi can.
In Fig. 4, we show a heatmap of the multi-object grasping success rate in simulation for our dataset replay experiments. Quantitatively, the results indicate that grasp poses from our dataset have potential for effective multi-object grasping. Most of the object-object grasping combinations result in a success rate above 80%, and four of them are at 100% (i.e., a 25 out of 25 success rate). The lowest observed success rate is 56%, occurring for object pairs involving the lemon, which is highly unstable on the tabletop and easily shifts upon contact with fingertips.
Overall, we achieve an average success rate of 82.7%, while for grasp poses generated by the diffusion model, the success rate drops to 65.8%. This decrease is expected due to the inherent noise in diffusion-based sampling [2]. See the supplementary website for more details.
Table II shows the result of our ablation study on the rotation representation choice for the hand pose. The results show that our rotation matrix representation results in the highest success rates, consistent with the findings of [4].
Euler Angle Axis Angle Quaternion 6D Ours
62.56% 60.25% 62.13% 64.25% 65.81%
TABLE II: Ablation study to compare rotation representations for hand poses. See Sections V-B and V-C for more details.
V-D Real-World Experiment Setup
[…]
We conduct two types of real-world evaluations, based on Dataset Replay (DR) and Learned Grasp (LG) execution. These evaluations are similar to those in the simulation experiments (see Section V-B), except that for LG, the diffusion model is conditioned on real-world point clouds.
To reduce the sim-to-real gap, we follow the approach of [50] to obtain real-world object point clouds. To achieve this, we use RGB images of the entire scene from different camera views. We use Nerfstudio [51] to compute COLMAP reconstructions [52] and Stable Normal [53] to generate normal maps. Then, we employ 2D Gaussian Splatting [54] to obtain the point clouds. To determine the camera’s pose relative to the scene, we use AprilTag markers [55]. Given that we assume a canonical pose for the object, it is straightforward to crop out the object’s point cloud.
For both dataset replay and learned grasp execution, we conduct five trials per object pair. Since scene reconstruction is time-consuming, we perform it once per object combination. We use our trained diffusion model, conditioned on points sampled from the resulting real-world object mesh, to generate grasp poses.
V-E Results from Real-World Experiments
We show an example successful trial in Fig. 1. See Fig. 5 for more example trials of SeqMultiGrasp. We report the results of our real-world experiments in Table III, which presents success rates for dataset replay (DR) and learned grasp execution (LG) across different object pairs. The average success rate for dataset replay is 64.4%, while the learned grasp execution achieves 56.7%.
For pinch-like grasps, the primary failure mode arises when the grasp pose does not achieve a stable force closure. This instability leads to the object slipping as the hand moves to grasp the second object. Fig. 6 shows a representative grasping failure mode during the first step.
For side grasps, the main failure mode is object sliding during finger squeezing. See the last row of Fig. 5 for an example of this failure mode, where the hand was unable to properly enclose the second object. Additionally, for the soda can, during execution, the thumb may collide with the can as the hand moves, preventing successful grasping.
slender cylinder soup can soda can
DR LG DR LG DR LG
plum 4/5 4/5 3/5 3/5 3/5 3/5
cube 5/5 3/5 2/5 3/5 2/5 1/5
duck 3/5 4/5 5/5 1/5 3/5 2/5
small cylinder 5/5 3/5 4/5 5/5 4/5 5/5
bunny 3/5 3/5 3/5 4/5 4/5 2/5
lemon 4/5 2/5 1/5 3/5 0/5 0/5
TABLE III: Real-world grasping results. “DR” refers to dataset replay, while “LG” refers to learned grasp execution using the diffusion model (see Section IV-E).
Figure 6: An example failure mode of SeqMultiGrasp during the pinch grasp stage (for the first object), where the cube slips out of control.
VI Limitations
While promising, SeqMultiGrasp has limitations that motivate future research directions. First, we only test sequentially grasping $N_{o}=2$ objects, but humans can often sequentially grasp more than two objects. Second, our dataset has relatively limited size and diversity compared to some in dexterous manipulation [2, 7] and we plan to massively increase the dataset. Third, SeqMultiGrasp relies on several heuristics and assumptions about the nature of the grasps, which may not hold in many situations. There is a high computation cost of processing COLMAP and 2D Gaussian Splatting. Finally, we only consider rigid objects, and leave sequential grasping of deformable objects to future work.
VII Conclusions
In this work, we propose SeqMultiGrasp, a method for sequentially grasping multiple objects with one hand. First, we generate feasible single-object grasps by optimizing an energy function and validating them in simulation. Next, we merge compatible single-object grasps to construct multi-object grasp configurations. Then, we train a generative model on this dataset to generate poses conditioned on observational point cloud data. We perform real-world experiments to perform sequential multi-object grasping. We hope this inspires future work on dexterous multi-object grasping.
ACKNOWLEDGMENTS
We thank Haozhe Lou, Zhiyuan Gao, Yiyang Ling, Chaoyi Pan, and Haoran Geng for their helpful technical advice and discussions. We are especially grateful to Yuyang Li for his invaluable help and support.

### v2

@@
 For real-world deployment, we train a diffusion model conditioned on point clouds to propose grasp poses, followed by a heuristic-based execution strategy.
-We test our system using $8\times 8$ object combinations in simulation and $6\times 3$ object combinations in real.
-Our diffusion-based grasp model obtains an average success rate of 65.8% over 1600 simulation trials and 56.7% over 90 real-world trials, suggesting that it is a promising approach for sequential multi-object grasping with multi-fingered hands.
+We test our system using $\times 8$ 8 object combinations in simulation and 6×36$\times$ 36 × 3 object combinations in real.
+Our diffusion-based grasp model obtains an average success rate of 65.8% over 1,600 simulation trials and 56.7% over 90 real-world trials, suggesting that it is a promising approach for sequential multi-object grasping with multi-fingered hands.
 Supplementary material is available on our project website: https://hesic73.github.io/SeqMultiGrasp.
@@
 Then, the hand rotates so that it can grasp (3) and lift (4) the can, without losing control of the plum.
-The contributions of this work are: (i) a novel framework, SeqMultiGrasp, for real-world sequential multi-object grasping; (ii) a new open-source dataset specifically designed to train sequential multi-object grasping methods; (iii) the first real-world robotic experiments demonstrating sequential multi-object grasping, where we conduct a total of 180 trials across 18 object pairs, with 90 trials replaying grasps from our dataset and 90 trials using grasp poses generated by our diffusion model.
+The contributions of this work are: (i) a novel synthesis pipeline for generating physically-feasible multi-object grasp configurations; (ii) an integrated system, SeqMultiGrasp, for real-world sequential grasping of multiple objects with a dexterous hand; (iii) the first real-world robotic experiments demonstrating sequential multi-object grasping, where we conduct a total of 180 trials across 18 object pairs, with 90 trials directly executing synthesized grasp poses and 90 trials using grasp poses generated by our diffusion model.
 II Related Work
@@
 To describe a single object, we can use $O$ and suppress the subscript when the distinction is not needed.
-Our objective is to generate a sequence of grasping actions $\mathcal{A}=\{a^{1},a^{2},\dots,a^{L}\}$ such that a robot sequentially grasps and lifts multiple objects while maintaining grasp stability for all objects in hand.
+Our objective is to generate a sequence of $L$ grasping actions, $\mathcal{A}=\{a^{1},a^{2},\dots,a^{L}\}$, such that a robot sequentially grasps and lifts multiple objects while maintaining grasp stability for all objects in hand.
 We assume a robot arm with a single multi-fingered hand.
@@
 IV-C Grasp Validation in Simulation
-While the optimization procedure from Section IV-B creates grasp candidates, we still require validation to ensure that they lead to successful grasps.
-Thus, we execute grasps in a GPU-accelerated physics simulator, ManiSkill [39], to validate them while filtering out unsuccessful candidates.
 […]
-Rotation Robustness: From [3], we evaluate grasp stability under six different gravity directions. A grasp is considered valid if it maintains force closure across all tested conditions.
+Rotation Robustness: From [3], we evaluate grasp stability under six axis-aligned gravity directions (±$\pm x,\pm y,\pm z$); a grasp is considered valid if the object remains in contact with the hand after 2.5 seconds of simulation across all tested conditions.
 Execution Feasibility: Adapted from [2], we verify whether the grasp can be successfully executed without causing collisions with the environment.
@@
 Similarly, the dashed red box shows the set of objects that are grasped second with a side grasp.
-See Tables I and III for the name of these objects in simulation and real, respectively.
+See Tables I and II for the name of these objects in simulation and real, respectively.
 The annotated numbers for simulation objects match those in Table I.
@@
 Specifically, we employ a simple squeeze-and-lift procedure.
-Given a hand configuration, we follow the strategy in Section IV-A to determine which fingers are used for each object.
+Although our execution procedure is not restricted to a specific finger-to-object assignment, we adopt the same strategy as in IV-A for consistency, as the diffusion model is biased toward the dataset’s distribution.
 During grasp execution, we first use CuRobo [45] for motion planning to move the end effector to a collision-free pose set at a position offset from the grasp pose.
@@
 V Experiments
-V-A Object Selection and Data Generation Evaluation
+V-A Object Selection and Grasp Synthesis Statistics
 For experiments in simulation, we select 8 objects for pinch-like grasping and 8 objects for side-grasping from YCB [46], ContactDB [47] and Omni6DPose [48].
@@
 3 (left).
-To evaluate our dataset generation pipeline, we report the synthesis success rates for pinch-grasping objects and side-grasping objects (Table I).
+To assess the efficiency of our single-object grasp pose synthesis method, we report the synthesis success rates for pinch-grasping objects and side-grasping objects (Table I).
 Due to the use of fewer hand links and our strict hybrid validation criteria, the success rates are lower compared to single-object grasping scenarios [3, 4, 2].
@@
 We conduct the following experiments in simulation:
-  • Dataset Replay (DR): We evaluate grasp poses sampled from our dataset.
-While each single-object grasp has already passed strict validation, failure may still occur after merging.
-This is mainly due to three factors: (1) Object placement shifts may lead to motion planning failures.
-(2) During execution, unintended interactions between the two objects or between the hand and objects may cause slippage or collisions.
-(3) A slight discrepancy exists between the validation and evaluation environments, as validation uses the PhysX GPU backend, while evaluation runs on the CPU.
+  • Synthesized Grasp (SG): We directly use grasp poses synthesized in the dataset for evaluation.
+Although each single-object grasp is strictly validated during dataset construction, failures may still occur.
+This is mainly due to: (1) object placement shifts leading to planning failures; (2) unintended interactions between the hand and objects or between the objects themselves; (3) a small simulation gap, as dataset validation is performed in batch on a GPU physics backend, while evaluation uses a CPU backend.
+This also serves as a baseline for the subsequent learned grasp evaluation.
   • Learned Grasp (LG): We sample grasp poses from our diffusion model conditioned on point clouds and execute them in simulation to measure the success rate.
-  • Ablation studies: We compare rotation representations for network prediction, including Euler angles, axis-angle, quaternion, 6D representation [49], and our chosen rotation matrix (orthogonalized via SVD).
-We test each of these representations over all $8\times 8$ object combinations using the LG experiment protocol.
-Figure 4: A heat map where each grid cell shows the success rate (from 25 trials) of sequentially grasping two objects in simulation for dataset replay experiments.
+Figure 4: A heat map where each grid cell shows the success rate (from 25 trials) of sequentially grasping two objects in simulation for synthesized grasp experiments.
 The first grasp is a pinch grasp (vertical axis) and the second grasp is a side grasp (horizontal axis).
@@
 In Fig.
-4, we show a heatmap of the multi-object grasping success rate in simulation for our dataset replay experiments.
+4, we show a heatmap of the multi-object grasping success rate in simulation for our synthesized grasp experiments.
 Quantitatively, the results indicate that grasp poses from our dataset have potential for effective multi-object grasping.
@@
 See the supplementary website for more details.
-Table II shows the result of our ablation study on the rotation representation choice for the hand pose.
-The results show that our rotation matrix representation results in the highest success rates, consistent with the findings of [4].
-Euler Angle Axis Angle Quaternion 6D Ours
-62.56% 60.25% 62.13% 64.25% 65.81%
-TABLE II: Ablation study to compare rotation representations for hand poses. See Sections V-B and V-C for more details.
 V-D Real-World Experiment Setup
 […]
-We conduct two types of real-world evaluations, based on Dataset Replay (DR) and Learned Grasp (LG) execution.
+We conduct two types of real-world evaluations, based on Synthesized Grasp (SG) and Learned Grasp (LG).
 These evaluations are similar to those in the simulation experiments (see Section V-B), except that for LG, the diffusion model is conditioned on real-world point clouds.
@@
 Given that we assume a canonical pose for the object, it is straightforward to crop out the object’s point cloud.
-For both dataset replay and learned grasp execution, we conduct five trials per object pair.
+For both synthesized and learned grasp evaluations, we conduct five trials per object pair.
 Since scene reconstruction is time-consuming, we perform it once per object combination.
@@
 5 for more example trials of SeqMultiGrasp.
-We report the results of our real-world experiments in Table III, which presents success rates for dataset replay (DR) and learned grasp execution (LG) across different object pairs.
-The average success rate for dataset replay is 64.4%, while the learned grasp execution achieves 56.7%.
+We report the results of our real-world experiments in Table II, which presents success rates for synthesized grasp (SG) and learned grasp (LG) across different object pairs.
+The average success rate for SG is 64.4%, while LG achieves 56.7%.
 For pinch-like grasps, the primary failure mode arises when the grasp pose does not achieve a stable force closure.
@@
 slender cylinder soup can soda can
-DR LG DR LG DR LG
+SG LG SG LG SG LG
 plum 4/5 4/5 3/5 3/5 3/5 3/5
@@
 lemon 4/5 2/5 1/5 3/5 0/5 0/5
-TABLE III: Real-world grasping results. “DR” refers to dataset replay, while “LG” refers to learned grasp execution using the diffusion model (see Section IV-E).
+TABLE II: Real-world grasping results. “SG” refers to synthesized grasp, while “LG” refers to learned grasp execution using the diffusion model (see Section IV-E).
 Figure 6: An example failure mode of SeqMultiGrasp during the pinch grasp stage (for the first object), where the cube slips out of control.
@@
 We are especially grateful to Yuyang Li for his invaluable help and support.
+We sincerely thank the IROS 2025 reviewers for their constructive feedback.